Pepper’s Products manufactures and sells two types of chew toys for pets—Squeaky and Silent. In May, Pepper’s Products had the following costs and revenues.

Pepper's Products
Income Statement
For the Month of May
Squeaky Silent Total
Sales revenue $ 138,000 $ 154,000 $ 292,000
Direct materials 19,000 22,000 41,000
Direct labor 80,000 20,000 100,000
Overhead costs
Administration 20,000
Production setup 45,000
Quality control 20,000
Distribution 21,000
Operating profit $ 45,000

Pepper’s Products currently uses labor costs to allocate all overhead but is considering implementing an activity-based costing system. After interviewing the sales and production staff, management decides to allocate administrative costs on the basis of direct labor costs but to use the following bases to allocate the remaining overhead.

Activity Level
Activity Cost Driver Squeaky Silent
Setting up Number of production runs 10 20
Performing quality control Number of inspections 40 40
Distribution Number of units shipped 80,000 130,000

Required:

a. Complete the income statement using the preceding activity bases.

c. Restate the income statement for Pepper's Products using direct labor costs as the only overhead allocation base.

Restate the income statement for Pepper's Products using direct labor costs as the only overhead allocation base. (Do not round intermediate calculations. Input all amounts as positive values.)

Account Squeaky Silent Total
Sales revenue $138,000 $154,000 $292,000
Direct materials 19,000 22,000 41,000
Direct labor 80,000 20,000 100,000
Overhead costs
Operating profit (loss)
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Pepper’s Products
Answer a A B C
D=B+C
E=A/D
Calculation of ABC rates for overhead R Cost Driver
Activity cost pool Cost Driver Cost Squeaky Silent Total ABC rates
Setting up Number of production runs        45,000.00                 10.00             20.00             30.00        1,500.00
Performing quality control Number of inspections        20,000.00                 40.00             40.00             80.00           250.00
Distribution Number of units shipped        21,000.00         80,000.00 130,000.00 210,000.00                0.10
Total Cost assigned        86,000.00
See E See B F=B*E See C G=C*E
Allocation table Squeaky Silent
Activity cost pool Cost Driver ABC Rates Activity used Cost assigned Activity used Cost assigned
Setting up Number of production runs          1,500.00                 10.00      15,000.00             20.00      30,000.00
Performing quality control Number of inspections              250.00                 40.00      10,000.00             40.00      10,000.00
Distribution Number of units shipped                  0.10         80,000.00        8,000.00 130,000.00      13,000.00
Total Cost assigned     33,000.00     53,000.00
Income Statement Squeaky Silent Total
Sales Revenue      138,000.00      154,000.00       292,000.00
Direct Materials        19,000.00        22,000.00         41,000.00
Direct Labor        80,000.00        20,000.00       100,000.00
Overhead costs
Production set up        15,000.00        30,000.00         45,000.00
Quality control        10,000.00        10,000.00         20,000.00
Distribution          8,000.00        13,000.00         21,000.00
Operating Profit (Loss)          6,000.00        59,000.00         65,000.00
Administration cost         20,000.00
Net Profit (Loss)         45,000.00
Calculation of single plant wide OH Rate Total Note
Direct Labor costs      100,000.00 H
Overhead costs        86,000.00 I
Predetermined OH Rate                  0.86 J=I/H
Overhead allocated to each product Squeaky Silent Note
Direct Labor costs        80,000.00        20,000.00 K
Predetermined OH Rate                  0.86                  0.86 See J
Overhead allocated to each unit        68,800.00        17,200.00 L=K*J
Income Statement Squeaky Silent Total
Sales Revenue      138,000.00      154,000.00       292,000.00
Direct Materials        19,000.00        22,000.00         41,000.00
Direct Labor        80,000.00        20,000.00       100,000.00
Overhead costs        68,800.00        17,200.00         86,000.00
Operating Profit (Loss)      (29,800.00)        94,800.00         65,000.00
Administration cost         20,000.00
Net Profit (Loss)         45,000.00